Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Yi Qiang
Log entry time 08:48:45 on December 05, 2008
Entry number 252128
This entry is a followup to: 251837
Followups:
keyword=Experiment Summary for Dec 3, Owl - Dec 4, Swing
Summary for Dec 3 owl - Dec 4 swing
Smooth data taking.
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/12/03 owl | 4.51 |
0.34 | 3.15 |
2008/12/03 day | 6.50 |
0.49 | 1.02 |
2008/12/03 swing | 4.18 |
0.15 | 3.67 |
2008/12/04 owl | 6.71 |
0.15 | 1.12 |
2008/12/04 day | 6.56 |
0.44 | 0.98 |
2008/12/04 swing | 5.34 |
2.06 | 0.60 |
total | 33.80 |
3.63 | 10.54 |
Main events/issues:
- Production afer 3:00am, 2008/12/03.
- TS problem after 22:00pm, 2008/12/04
Plans for the next few days:
- Change HRS polarity during Friday Day shift.
- Swape TS module in the same access of polarity change.